Exercise 4-1 (Video) Saddle Ine. has two types of handbags: standard and custom. The controller has decided to use a plantwide overhead rate based on direct labor costs. The president has heard of activity-based costing and wants to see how thie results would dmer 년 this system were used. Two activty cost pools were developed: machining and machine setup. Presented below is iformation related to the companys operations. Standard Custom Direct labor Machine hours Setup hours 50,000 200,000 1,200 420 1,500 Total estimated overhead costs are $285,900. Overhead cost allocated to the machining activity cost pool is $189,000, and $96,900 is allocated to the machine setup activity cost pool E] Your answers correct. Compute the overheed rate using the traditional (plantwide) approach. (Round answer to 2 decimal placeso.g. 1225) 190.60 % of direct labor cost 回Your rowers correct. 区] Compute the overhead rates using the activity-based costing approach Machining Machine 70 per machine hour 190 per setup hour
回Your answer is correct. Compute the overhead rates using the activity-based costing approach Machining per machine hour 190l per setup hour MILAR EXERCINE Your answer is incerrect. Try agai. Determine the derence in allocation between the two approaches Standard Standard Custom Click if you would like to Show Work for this questions een hoWork
Just answer the ones i got wrong please :) Thanks!
Answer #1

Overhead Allocation as per traditional costing:

Standard: 50,000*190.60% = $95,300

Custom: 100,000*190.60% = $190,600

Using Activity Based Costing:

Standard:

Machining 1500*70 + Set Up 90*190 = $122,100

Custom: 1,200*70 + 420*190 = $163,800